OpenStudy (anonymous):

Hello, I am trying to help my son with his algebra! but i am just having hardest time! can someone please help me?

hartnn (hartnn):

Hi @janice1989 \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{E}}\color{#e6e600}{\text{L}}\color{green}{\text{C}}\color{blue}{\text{O}}\color{purple}{\text{M}}\color{purple}{\text{E}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{P}}\color{purple}{\text{E}}\color{red}{\text{N}}\color{orange}{\text{S}}\color{#e6e600}{\text{T}}\color{green}{\text{U}}\color{blue}{\text{D}}\color{purple}{\text{Y}}\color{purple}{\text{ }}\color{red}{\text{:}}\color{orange}{\text{)}}\color{#e6e600}{\text{}}\end{array} \]

hartnn (hartnn):

how can we help you ? do you have any specific qustion ?

OpenStudy (anonymous):

Yes can you give me a min? I am not sure how to really work this nad help my son, he has special needs! so he has hard time understanding it and im just stuck! and the teachers dont help at all! im just os stressed out!

hartnn (hartnn):

don't worry, i will try my best to help you :)

hartnn (hartnn):

you just post the question you have difficulty with.

OpenStudy (anonymous):

Okay, Sorry it took so long to reply he got one asnwer right hold on whille i type the ones we need help with! What is your name?

hartnn (hartnn):

my name is hartnn, janice :) you can call me that

OpenStudy (anonymous):

Okay!

OpenStudy (anonymous):

The expression 5x^2+x-10 has ___ terms and a constand of ____

OpenStudy (anonymous):

Oh i forgot answers

OpenStudy (anonymous):

3,-10 3,10 6,-10 6,10 My son said B

hartnn (hartnn):

3 terms are correct! 5x^2 -> one term +x ->2nd term -10 -> 3rd term but the constant is the term with no ''x" in it, which is ... ?

OpenStudy (anonymous):

-10 he said

hartnn (hartnn):

and he is correct! :)

OpenStudy (anonymous):

So the answer would be A than correct?

hartnn (hartnn):

yes, A would be correct.

OpenStudy (anonymous):

He is smiling ear to ear he got it right! you can explain it a lot better than i could, there is some more if you dont mind helping with !

hartnn (hartnn):

i don't mind, at all :)

OpenStudy (anonymous):

Janice, thanks for using OpenStudy. Do tell other parents about us. Hartnn is one of our senior, most skilled and patient OpenStudiers! There are many like him on the site to help. They do it because they love Math and helping young learners.

OpenStudy (anonymous):

the expression 9x^2-5x-7 has __ terms and a constant of ___ 3,-7 5,-7 3,7 -5,7 He said there is 5 terms and a constant of -7 since its right after the x. so he said it would be B correct?

OpenStudy (preetha):

Good work Hartnn!

OpenStudy (anonymous):

I will for sure tell other parents about it! Ive been stuck on this for such a long time, And someone told me about it, this is an amazing website, Do you also do Geometry?? My daughter is doing geometry and also has special needs she has Aspergers and is having such a hard time with it.. and she is just stressed and all!

hartnn (hartnn):

thank you for the appreciations, yes, i do geometry too :) lets count the terms in , 9x^2-5x-7 9x^2 -5x -7 how many ?

OpenStudy (anonymous):

Than after we finish this up maybe later on you could help my daughter? And he said 3 :)

OpenStudy (anonymous):

so instead he said would it be A than?

hartnn (hartnn):

A is correct :)

OpenStudy (anonymous):

Woohoo you are amazing!!!

OpenStudy (anonymous):

In 6x^4-3y+5x^4-10 what are the like terms in the expression choose all that apply.. 6x^4 -3y 5x^4 -10 he said all of them

hartnn (hartnn):

you will find many amazing users here on OpenStudy, willing to help you as much as i am :)

hartnn (hartnn):

like terms are those which have same exponent of 'x' like x^3, -3x^3, 8x^3 .... (one set of like terms) x,4x,9x,-x,.... and constants like 4,8,0,-1,....

hartnn (hartnn):

in 6x^4-3y+5x^4-10 which terms have same exponent of x ?

OpenStudy (anonymous):

so would it be A & C

hartnn (hartnn):

does your boy say/agree with A and C too ?

OpenStudy (anonymous):

Yeah he said that, i am just typing for him because he is little bit slow on typing! :D

hartnn (hartnn):

tell him he is absolutely correct :)

OpenStudy (anonymous):

Okay :D hold on he is working on next one to see if he can get it, who would have known a stranger online could help so much more than a teacher!

hartnn (hartnn):

now you know :)

OpenStudy (anonymous):

Simplify : 7x^2+2x^2 7x^2 9x^2 14x^4 9x^4 He said he is not sure what to do can you explain it step by step? He said it would either be A or B since the top number are both 2. or do you add them together?

OpenStudy (anonymous):

He said would it be B? since 7+2 = 9??

hartnn (hartnn):

the top number (here 2) is called the exponent. does he know what an exponent is ? and yes, B is correct, because 7x^2+2x^2 = (7+2)x^2 = 9x^2

OpenStudy (anonymous):

He said hes heard of it, but isnt really sure what it is.

hartnn (hartnn):

i'll explain, x^n, x is called the 'base' and n is called the exponent. x^n implies , multiplying 'x' to itself, 'n' times. we take an example : 2^4 will imply multiplying 2 4 times = 2*2*2*2 = 16 similarly, x^2 will imply multiplying x with itself twice, x*x got this ?

OpenStudy (anonymous):

he said oh that is what exponent is i know now! lol. Thank you, hes working on the next one :)

OpenStudy (anonymous):

Subtract the second expression from the first (5x^2-3x+5)-(3x^2-4x-7) 2x^2+x+12 2x^2-7x+12 8x^2+x+12 2x^2+x-2 He said would it be B?

hartnn (hartnn):

why B , how did he get co-efficients like 2,-7 and 12 ?

OpenStudy (anonymous):

He says he dont know, i think he was guessing. He does that a lot. Im sorry!

hartnn (hartnn):

(5x^2-3x+5)-(3x^2-4x-7) = 5x^2 -3x +5 -3x^2 - (-4x) - (-7) {by distributing the negative} = 5x^2 -3x +5 -3x^2 +4x +7 {because negative times negative becomes positive!} then we identify like terms. but is this much clear ?

OpenStudy (anonymous):

he said for like terms would it be 5x^2 -3x -3x^2 and 4x?

hartnn (hartnn):

for like terms , the exponent of x should be same, for which terms exponent of x is 2 ? for which terms exponent of x is 1 ?

hartnn (hartnn):

note x means x^1 , both mean the same

OpenStudy (anonymous):

he said he is confused. he said he dont get that,

hartnn (hartnn):

ok, the 5x^2 and -3x^2 has exponent of 2 so these are like terms and we can combine them in this way, 5x^2-3x^2 = (5-3)x^2 = 2x^2 got this?

OpenStudy (anonymous):

got it he said

hartnn (hartnn):

cool, from this : 5x^2 -3x +5 -3x^2 +4x +7 the other like terms are -3x and +4x (because they both contains just x with it, not x^2 or x^3 .....) so we can combine them like : -3x+4x = (-3+4) x = ... ?

OpenStudy (anonymous):

he said uh he thinks you can combine them not 100% sure

hartnn (hartnn):

yes, they can be combined because they both have same exponent (=1 here) so, -3x+4x = (-3+4) x = -x lastly, what remains are just constants, 7+5 , which =12

OpenStudy (anonymous):

ok he got this,

OpenStudy (anonymous):

he understands whats next step?

hartnn (hartnn):

so, finally, putting the results together, 2x^2-x+12

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

distribute and simplify 7(2x-4)+2(3x+1) 20x+26 20x-26 14x-28 6x+2 He asked what is it to distribute???

hartnn (hartnn):

Distribution Property : a(b+c) = ab+ac  ask if you need an explanation in words....